# REAL-Prover: Retrieval Augmented Lean Prover for Mathematical Reasoning

### v2

@@ Abstract
Nowadays, formal theorem provers have made monumental progress on high-school and competition level mathematics, but few of them generalize to more advanced mathematics. In this paper, we present REAL-Prover, a new open-source stepwise theorem-prover for Lean 4 to push this boundary. This prover, based on our fine-tuned large language model (REAL-Prover-v1) and integrated with a retrieval system (Leansearch-PS), notably boosts performance on solving college-level mathematics problems. To train REAL-Prover-v1, we developed HERALD-AF, a data extraction pipeline that converts natural language math problems into formal statements, and a new open-source Lean 4 interactive environment (Jixia-interactive) to facilitate synthesis data collection. In our experiments, our prover using only supervised fine-tune achieves competitive results with a 23.7% success rate (Pass@64) on the ProofNet dataset—comparable to state-of-the-art (SOTA) models. To further evaluate our approach, we introduce FATE-M, a new benchmark focused on algebraic problems, where our prover achieves a SOTA success rate of 56.7% (Pass@64).
1 Peking University  2 Renmin University of China  3 National University of Singapore  4 Ubiquant
5 Beijing International Center for Mathematical Research and the New Cornerstone Science Laboratory, Peking University
6 Center for Machine Learning Research, Peking University
7 Center for Intelligent Computing, Great Bay Institute for Advanced Study, Great Bay University
dongbin@math.pku.edu.cn
Code: https://github.com/frenzymath/REAL-Prover
@@ 1 Introduction
The formalization of mathematics has long been a great challenge at the intersection of programming and mathematical logic. Over the past decades, theorem provers have made steady progress. In particular, proof assistants have been used to formalize complex theorems such as the Four-Color Theorem fourcolourthm and the Feit-Thompson Odd-Order Theorem oddthm . These achievements underscore the potential of formal methods to resolve research-level proofs.
@@
Modern Interactive Theorem Provers (ITPs) like Lean lean and Coq coq have significantly improved usability and automation, lowering the barrier for adoption. Lean, for example, provides a user-friendly environment and a large community-driven library of mathematics (Mathlib) The_mathlib_Community_2020 . In early 2025, Lean’s Mathlib had grown to encompass more than 210,000 formally proven theorems in various domains. Meanwhile, utilities like LeanSearch gao2025semanticsearchenginemathlib4 also emerge to help users navigate easily in these numerous domains and facts, saving their efforts on formal proof drafting.
Recent advances in large language models (LLMs) have significantly advanced automated theorem proving, successfully generating formal proofs for high-school-level math competitions polu2020generativelanguagemodelingautomated ; jiang2023draftsketchproveguiding ; zhao2023decomposingenigmasubgoalbaseddemonstration ; wang2023legoproverneuraltheoremproving ; xin2024deepseekprover ; wu2024internlm2 ; xin2024deepseekproverv15harnessingproofassistant ; xin2025bfsproverscalablebestfirsttree ; ren2025deepseekproverv2advancingformalmathematical . However, extending these approaches to undergraduate and graduate mathematics remains challenging due to data sparsity, domain-specific complexities, and the necessity of accurately recalling extensive lemma libraries, where even minor errors can derail purely neural proof generation.
To overcome these challenges, we propose a stepwise tactic synthesis approach in which each invocation of the language model is conditioned on the current proof state and enhanced by retrieving pertinent theorems from Mathlib. First, we introduce LeanSearch-PS, a semantic premise selection engine inspired by LeanSearch gao2025semanticsearchenginemathlib4 , designed to efficiently identify the $k$-most relevant theorems for any given proof state. Specifically, at each tactic step, we extract the current proof state from the Lean compiler and then query LeanSearch-PS to obtain candidate theorems. Second, we introduce REtrieval Augmented Lean Prover, named REAL-Prover, an integrated framework which combines the proof state information and these retrieved premises as inputs for the tactic-generation model. Specifically, it consists of an expert iteration framework and a translator pipeline. The former starts by imitating a small dataset derived from expert-guided proofs, then iteratively refines the prover through self-play interactions between the prover and compiler. The latter, HERALD-AF, is an extension of the HERALD translator model to be a whole translation pipeline gao2025heraldnaturallanguageannotated , which generates and validates formal statements automatically from mathematical textbooks.
Furthermore, to support efficient interaction between prover and compiler in stepwise proving, we also developed Jixia-interactive, a Lean-based interactive environment. Specifically, the environment tool works within an all-Lean, CLI-driven scope with transparent environment and version management. It processes Lean tactic interactions internally, while collecting and presenting errors/warnings as the exploration or edition goes deeper. Compared to previous static-analysis tools like REPL, JixiajixiaGitHub and LeanDojoyang2023leandojotheoremprovingretrievalaugmented , Jixia-interactive is more usable, stable and efficient.
We finally propose a benchmark Formal Algebra Theorem Evaluation-Medium, named FATE-M, which is designed for formal reasoning in graduate-level abstract algebra. It enriches existing benchmarks emphasizing advanced algebraic structures (e.g., groups, rings, fields), making it an ideal testbed for exploring complex formal proofs. We evaluate our REAL-Prover on FATE-M, achieving a state-of-the-art accuracy of 56.7%, and further demonstrate its effectiveness on general benchmarks miniF2F miniF2F and ProofNet azerbayev2023proofnetautoformalizingformallyproving , where it achieves competitive results.
We summarize our contribution as follows.
  • Retrieval-Augmented Proof Search. We propose REAL-Prover, a stepwise theorem proving system integrated with LeanSearch-PS, a semantic premise-selection engine. This approach significantly enhances prover effectiveness on graduate-level formal mathematics.
  • Automated Data Generation Pipeline. We develop HERALD-AF, a novel pipeline for translating informal mathematical statements into formal Lean statements. This system facilitates large-scale auto-formalization, substantially enriching the available training dataset.
  • Lean Interactive Environment. We introduce Jixia-interactive, an advanced interactive platform for Lean 4, designed to provide efficient, stable, and user-friendly interactions between the prover and compiler throughout both training and inference phases.
  • Specialized Benchmark for Abstract Algebra. We present FATE-M (Formal Algebra Theorem Evaluation–Medium), a specialized benchmark comprising paired informal-formal problems in graduate-level abstract algebra. FATE-M complements existing benchmarks such as miniF2F miniF2F and ProofNet azerbayev2023proofnetautoformalizingformallyproving , establishing an effective framework for rigorous formal reasoning assessment.
@@ 2 Method
To construct a high-performance automated prover, we propose REtrieval-Augmented Lean Prover (REAL-Prover), a stepwise proof system for Lean 4. To supply REAL-Prover with a high-quality training corpus, we first introduce Hierarchy and Retrieval-based Translated Lean Dataset Auto-formalization (HERALD-AF) in Section 2.1, a pipeline that translates informal mathematical problems into formal Lean 4 statements. HERALD-AF converts informal mathematical problems into formal Lean 4 statements, yielding a large corpus of candidate theorems. Since these statements still lack certified proofs, in Section 2.2, we build a complementary pipeline—again driven by REAL-Prover—that synthesizes and automatically verifies the required proofs, completing the training loop.
@@ 2.1 HERALD-AF
Our formal statement generation framework builds upon the Hierarchy and Retrieval-based Translated Lean Dataset (HERALD) gao2025heraldnaturallanguageannotated , which establishes a protocol translating informal statements into formal statements. Figure 1(a) schematizes HERALD-AF pipeline.
@@
Once the informal statements are obtained, we apply HERALD-AF to translate them into formal statements. This translation pipeline comprises three key stages: Auto-Formalization, Auto-Informalization, and LLM Judgement. In the Auto-Formalization stage, an open-source model Herald-translator gao2025heraldnaturallanguageannotated is used to generate multiple candidate formal statements from each informal input. During Auto-Informalization, these formal candidates are translated back into natural language using DeepSeek-V3 deepseekai2025deepseekv3technicalreport . Finally, in the LLM Judgement stage, a general LLM evaluates whether each back-translated statement faithfully reflects the original input. Only those formal statements that pass this consistency check are retained.
Figure 2: Overview of REAL-Prover. This figure illustrates three main processes: (a) to (b), the prover passes several tactics to the Lean 4 interactive environment (Jixia-interactive) and receives corresponding scored state nodes. (b) to (c), the prover selects a state node based on its score and checks whether the proof is complete. (c) to (a), the prover using LLM with retrieval information from LeanSearch-PS generate new tactics.
@@ 2.2 REAL-Prover
From the HERALD-AF pipeline, we obtained a collection of formal statements without corresponding proofs. To generate complete proofs and achieve the second goal, we introduce the REAL-Prover system, illustrated in Figure 2. We developed a new Lean 4 interactive environment, Jixia-interactive, which receives Lean tactics and updates the proof state. For each formal statement, Jixia-interactive initializes the proof state. REAL-Prover constructs a search tree to explore possible proof paths. At each node, a LLM generates multiple candidate tactics. Jixia-interactive then applies these tactics, discards invalid ones, and advances to new proof states. Similar to BFS-Prover xin2025bfsproverscalablebestfirsttree , each node is scored using the scoring function:$\frac{\sum_{t=0}^{L-1}\log p(a_{t}|s_{t})}{L^{\alpha}}$, where the $s_{t}$ represents the proof state and $a_{t}$ is the tactic applied at step t. $p(a_{t}|s_{t})$ is the model’s predicted probability of applying tactic $a_{t}$ at state $s_{t}$. $L$ is the total length of the path from the root to the current state $s_{L}$ and $\alpha$ is a hyper-parameter in $[0,1]$. We adopt a best-first search strategy, selecting the node with the highest score to expand further by generating new tactics. This iterative process continues until either the proof is completed or a predefined search budget is exhausted. Once completed, the full proof can be reconstructed from the search tree. To enhance our prover’s performace, we add retrieval information from Mathlib using LeanSearch premise selection (LeanSearch-PS), which extends the work of LeanSearch gao2025semanticsearchenginemathlib4 . We will discuss the details of LeanSearch-PS.
@@ 2.2.1 LeanSearch-PS
To train the embedding model, we construct a dataset of $(s,t_{\text{pos}})$ pairs. These pairs are extracted from Mathlib using the Jixia tool jixiaGitHub . In this context, $s$ refers to a Lean proof state, and $t_{\text{pos}}$ refers to a theorem that was successfully applied at that step. The training process follows a two-stage approach, inspired by transfer learning techniques commonly used to improve text retrieval SFRAIResearch2024 :
  • Initial Training: The model was initially trained on $(s,t_{\text{pos}})$ pairs using a contrastive loss function: $L(x_{1},x_{2})=yd(x_{1},x_{2})^{2}+(1-y)\max(m-d(x_{1},x_{2}),0)^{2}$. where the $d(x_{1},x_{2})$ denotes a distance metric between embeddings, and $y=1$ if $x_{1}$ and $x_{2}$ are similar, and $y=0$ otherwise. The $m$ is hyperparameter that defines the minimum distance between dissimilar pairs. In our data, we treat state as $x_{1}$ and $t_{\text{pos}}$ as $x_{2}$. We adopt the in-batch negatives strategy, where other premises within the same training batch that are not associated with the current state are treated as negative examples for contrastive learning of premise relevance.
  • Hard Negative Enhanced Training: To improve fine-grained discrimination, the initial model was used to mine hard negatives—challenging incorrect premises for each state. This process produces triplets of the form $(s,t_{\text{pos}},t_{\text{hard\_neg}})$, where $t_{\text{hard\_neg}}$ refers to hard negative premise. We then train the model using a triplet loss function: $L(x,p,n)=\max(d(x,p)-d(x,n)+m,0)$, where $x$ is state, $p$ is $t_{\text{pos}}$ and $n$ is $t_{\text{hard\_neg}}$. After this stage training, we obtain our final embedding model.
@@ 2.2.2 Expert Iteration
Another key component of REAL-Prover is the prover model within the tactic generator. To train a strong prover model, we combine expert iteration to collect synthetic data, as shown in Figure 1(b). We begin by fine-tuning a prover model using open-source datasets. This initial model is then employed within REAL-Prover to solve formal statements produced by HERALD-AF, generating additional high-quality data. The newly collected data is incorporated into the supervised fine-tuning dataset to train an improved model. This refined model is then used to tackle challenging formal statements that the previous model could not solve. We update our model continuously in this iteration process.
@@ 3.1 Implementation Detail
We perform supervised fine-tuning on the base models Qwen2.5-Math-7B yang2024qwen2 using a learning rate of $5\times 10^{-5}$ with a cosine decay scheduler and a maximum context length of 8192 tokens. Our training data comprises the following sources, totaling 210,420 state-tactic pairs:
  1. Post-processed formal proofs collected via expert iteration. Each proof is parsed into state-tactic pairs using Jixia, and the corresponding prompts are enriched with natural language descriptions, the proof context, the current proof state, and retrieval information from LeanSearch-PS. The dataset comprises 15,071 algebra problems from undergraduate and graduate mathematics texts, and 19,633 NuminaMath problems numina_math_datasets , resulting in 25,818 and 30,589 state-tactic pairs, respectively, after post-processing.
@@
  4. The augmented Lean-Workbook dataset ying2024leanworkbooklargescalelean . We enhance this dataset by adding retrieval information from LeanSearch-PS to each example, while keeping the original structure intact. The resulting dataset comprises 25,214 state-tactic pairs.
@@
We develop two dense retrievers from E5-mistral-7b-instruct wang2023improving with Tevatron Gao2022TevatronAE training framework. For both initial training and hard negative enhanced training, we use LoRA to accelerate training, and set the learning rate as $2\times 10^{-5}$, the batch size as 128, the number of epoch as 1, and other hyperparameters as their default values. Queries are informal statements with a maximum length of 128 and passages are formal theorems with a maximum length of 256. For the hard negative examples, we first embed all statements and theorems with the initial trained embedding model, and then randomly select one passage from the top 30 to top 100 most similar ones for each query as its hard negative premise.
@@
In our experiments, considering the overall computational budget, we set the number of passes and samples per step to 64. We set our LLM’s temperature to 1.5. Additionally, we set the hyperparameter $\alpha$ of the score function to 0.5, consistent with the configuration used in BFS-Prover xin2025bfsproverscalablebestfirsttree

### v3

@@ Abstract
Nowadays, formal theorem provers have made monumental progress on high-school and competition level mathematics, but few of them generalize to more advanced mathematics. In this paper, we present REAL-Prover, a new open-source stepwise theorem-prover for Lean 4 to push this boundary. This prover, based on our fine-tuned large language model (REAL-Prover-v1) and integrated with a retrieval system (Leansearch-PS), notably boosts performance on solving college-level mathematics problems. To train REAL-Prover-v1, we developed HERALD-AF, a data extraction pipeline that converts natural language math problems into formal statements, and a new open-source Lean 4 interactive environment (Jixia-interactive) to facilitate synthesis data collection. In our experiments, our prover using only supervised fine-tuning achieves competitive results with a 23.7% success rate (Pass@64) on the ProofNet dataset—comparable to state-of-the-art (SOTA) models. To further evaluate our approach, we introduce FATE-M, a new benchmark focused on algebraic problems, where our prover achieves a SOTA success rate of 56.7% (Pass@64).
keywords:
Neural Theorem Proving, Large Language Models, Formal Reasoning, Proof Search
[…]
68T15, 03B35, 68T20
@@ 1 Introduction
The formalization of mathematics has long been a great challenge at the intersection of programming and mathematical logic. Over the past decades, theorem provers have made steady progress. In particular, proof assistants have been used to formalize complex theorems such as the Four-Color Theorem [fourcolourthm] and the Feit-Thompson Odd-Order Theorem [oddthm]. These achievements underscore the potential of formal methods to resolve research-level proofs.
@@
Modern Interactive Theorem Provers (ITPs) like Lean [lean] and Coq [coq] have significantly improved usability and automation, lowering the barrier for adoption. Lean, for example, provides a user-friendly environment and a large community-driven library of mathematics (Mathlib) [The_mathlib_Community_2020]. In early 2025, Lean’s Mathlib had grown to encompass more than 210,000 formally proven theorems in various domains. Meanwhile, utilities like LeanSearch [gao2025semanticsearchenginemathlib4] also emerge to help users navigate easily in these numerous domains and facts, saving their efforts on formal proof drafting.
@@
Recent advances in large language models (LLMs) have significantly advanced automated theorem proving, successfully generating formal proofs for high-school-level math competitions [polu2020generativelanguagemodelingautomated, jiang2023draftsketchproveguiding, zhao2023decomposingenigmasubgoalbaseddemonstration, wang2023legoproverneuraltheoremproving, xin2024deepseekprover, wu2024internlm2, xin2024deepseekproverv15harnessingproofassistant, xin2025bfsproverscalablebestfirsttree, ren2025deepseekproverv2advancingformalmathematical]. Nonetheless, extending these approaches to undergraduate and graduate-level mathematics remains a significant challenge due to data sparsity, the intricacies of domain-specific reasoning, and the requirement for precise retrieval and application of extensive lemma libraries, where even minor inaccuracies can compromise the correctness of purely neural proof generation.
To address these challenges, we propose REAL-Prover (REtrieval-Augmented Lean Prover), a retrieval-augmented and stepwise framework for automated theorem proving in Lean. Each model invocation is conditioned on the current proof state and enhanced through the retrieval of semantically relevant theorems from Mathlib. The framework comprises three key components. LeanSearch-PS performs semantic premise selection through a two-stage training strategy that improves both retrieval precision and generalization. HERALD-AF extends the HERALD pipeline to systematically translate informal mathematical statements into verified Lean expressions, enabling scalable data generation. Jixia-Interactive provides a Lean 4 interactive environment that support efficient interaction between prover and compiler in stepwise proving. For evaluation, we construct FATE-M (Formal Algebra Theorem Evaluation–Medium), a benchmark designed for graduate-level reasoning in algebraic domains including groups, rings, and fields. REAL-Prover achieves a state-of-the-art accuracy of 56.7% on FATE-M and demonstrates competitive performance on miniF2F and ProofNet, confirming its effectiveness in complex formal reasoning.
1.1 Related works
Automated Theorem Proving
Automated Theorem Proving in formal languages, particularly within the framework of ITPs, is a rapidly advancing research area in artificial intelligence. LLM-based provers can be broadly categorized by their proof generation strategies: whole-proof generation that aims to produce complete proof in a single pass, and proof-step generation that iteratively refines proofs through ITPs.
The whole-proof generation paradigm focuses on end-to-end synthesis of formal proof scripts from theorem statements. Early attempts like DeepSeek-Prover-V1 [xin2024deepseek] demonstrated the feasibility of this approach in Lean 4. Subsequent improvements in DeepSeek-Prover-V1.5 [xin2024deepseekproverv15harnessingproofassistant] introduced hybrid techniques combining Reinforcement Learning from Proof Assistant Feedback (RLPAF) with Group Relative Policy Optimization [shao2024deepseekmath], along with a ‘truncate-and-resume’ mechanism that allows error correction through partial stepwise verification. Goedel-Prover [lin2025goedel] addressed data scarcity through auto-formalization, first converting natural language mathematics into formal statements before generating proofs through iterative training. While these approaches benefit from large-scale supervised fine-tuning and reduced ITP interaction, they struggle with error accumulation in long logical chains and implicit state inference.
This limitation has driven the development of interactive proof-step generation approaches that leverage ITPs feedback during the proving process. Systems like BFS-Prover [xin2025bfsproverscalablebestfirsttree] employ expert iteration with Best-First Search, using Lean compiler feedback to guide tactic generation. HunyuanProver [li2024hunyuanprover] enhances search efficiency through specialized critic models that evaluate intermediate proof states. The LeanDojo framework introduces ReProver [yang2023leandojotheoremprovingretrievalaugmented], which adopts a distinct retrieval-based approach using Dense Passage Retriever [karpukhin2020dense] to directly predict relevant premises. LeanAgent [kumarappan2024leanagent] extends this paradigm through lifelong learning, maintaining a dynamic knowledge base to progressively master theorem proving across multiple repositories. These interactive methods typically combine expert iteration, enabling real-time adaptation to proof state changes while managing computational complexity through guided tree search.
Formal Statement Benchmark
A fundamental aspect of developing and evaluating theorem provers is the utilization of formal statement benchmarks. Prominent examples include miniF2F [miniF2F], which offers a curated set of Olympiad-level formalization problems, and ProofNet [azerbayev2023proofnetautoformalizingformallyproving], which targets undergraduate-level mathematical reasoning. These benchmarks serve as standardized tools to assess the capabilities of theorem provers in generating correct and complete proofs. As the proficiency of provers advances, there is an increasing need for more challenging benchmarks, such as graduate-level or research-oriented problems, to rigorously evaluate their reasoning depth, adaptability, and ability to handle complex or previously unseen formalizations.
1.2 Our Contributions
Our main contributions are summarized as follows:
  • Retrieval-Augmented Proof System. We present a unified proof system that integrates data collection, training, and inference. The system includes three main components. HERALD-AF provides a pipeline for translating informal mathematical statements into formal Lean expressions. REAL-Prover implements a stepwise theorem-proving framework with LeanSearch-PS for semantic premise selection and employs an expert-iteration training paradigm. Jixia-Interactive facilitates efficient and stable interactions with the Lean 4 prover during both training and inference. The integration of these components significantly enhances proof performance on graduate-level formal mathematics.
  • State-Tactic Pair Dataset. We release a comprehensive dataset consisting of 55k state–tactic pairs generated through our expert iteration process. Each data instance encapsulates a formal proof state, its surrounding context, and the corresponding tactic sequence, providing high-quality supervision for both proof search optimization and policy learning.
  • Specialized Benchmark for Abstract Algebra. We present FATE-M (Formal Algebra Theorem Evaluation–Medium), a specialized benchmark comprising paired informal-formal problems in graduate-level abstract algebra. FATE-M complements existing benchmarks such as miniF2F [miniF2F] and ProofNet [azerbayev2023proofnetautoformalizingformallyproving], establishing an effective framework for rigorous formal reasoning assessment.
@@ 2 Method
To construct a high-performance automated prover, we propose REtrieval-Augmented Lean Prover (REAL-Prover), a stepwise proof system for Lean 4. To supply REAL-Prover with a high-quality training corpus, we first introduce Hierarchy and Retrieval-based Translated Lean Dataset Auto-formalization (HERALD-AF) in Section˜2.1, a pipeline that translates informal mathematical problems into formal Lean 4 statements. HERALD-AF converts informal mathematical problems into formal Lean 4 statements, yielding a large corpus of candidate theorems. Since these statements still lack certified proofs, in Section˜2.2, we build a complementary pipeline—again driven by REAL-Prover—that synthesizes and automatically verifies the required proofs, completing the training loop.
@@ 2.1 HERALD-AF
Our formal statement generation framework builds upon the Hierarchy and Retrieval-based Translated Lean Dataset (HERALD) [gao2025heraldnaturallanguageannotated], which establishes a protocol translating informal statements into formal statements. Figure˜1(a) schematizes HERALD-AF pipeline.
@@
Once the informal statements are obtained, we apply HERALD-AF to translate them into formal statements. This translation pipeline comprises three key stages: Auto-Formalization, Auto-Informalization, and LLM Judgement. In the Auto-Formalization stage, an open-source model Herald-translator [gao2025heraldnaturallanguageannotated] is used to generate multiple candidate formal statements from each informal input. During Auto-Informalization, these formal candidates are translated back into natural language using DeepSeek-V3 [deepseekai2025deepseekv3technicalreport]. Finally, in the LLM Judgement stage, a general LLM evaluates whether each back-translated statement faithfully reflects the original input. Only those formal statements that pass this consistency check are retained.
Figure 2: Overview of REAL-Prover. This figure illustrates three main processes: From (a) to (b), the prover passes several tactics to the Lean 4 interactive environment (Jixia-interactive) and receives corresponding scored state nodes. From (b) to (c), the prover selects a state node based on its score and checks whether the proof is complete. From (c) to (a), the prover using LLM with retrieval information from LeanSearch-PS generate new tactics.
@@ 2.2 REAL-Prover
[…]
@@ 2.2.1 LeanSearch-PS
To train the embedding model, we construct a dataset of $(s,t_{\text{pos}})$ pairs extracted from Mathlib using the Jixia tool [jixiaGitHub]. Here, $s$ denotes a Lean proof state, and $t_{\text{pos}}$ denotes the theorem successfully applied at that step. The training process adopts a two-stage framework inspired by transfer learning strategies commonly used in text retrieval [SFRAIResearch2024]. In both stages, we employ the InfoNCE loss[oord2018representation] defined as
$L(q,k_{i})=-\log\frac{\exp(q\cdot k_{0}/\tau)}{\sum_{i=0}^{n}\exp(q\cdot k_{i}/\tau)},$ (1)
where $k_{0}$ is the embedding of positive example, $k_{1},\ldots,k_{n}$ are the embedding of negative examples, and $\tau$ is a temperature hyperparameter. In the subsequent description, we will omit explicit mention of the embedding for brevity.
The training process consists of the following two stages.
  • Initial Training: We treat the proof state $s$ as the query $q$ and its corresponding theorem $t_{\text{pos}}$ as the positive example $k_{0}$. The in-batch negatives strategy is used, where other premises in the same batch serve as negative examples. This enables contrastive learning of premise relevance across proof states.
  • Hard Negative Enhanced Training: To refine fine-grained discrimination, we leverage the initially trained model to mine hard negatives—challenging but incorrect premises for each state. This yields triplets $(s,t_{\text{pos}},t_{\text{hard\_neg}})$, where $t_{\text{hard\_neg}}$ denotes a mined hard negative premise. During this stage, we use $s$ as $q$, $t_{\text{pos}}$ as $k_{0}$, and $t_{\text{hard\_neg}}$ as $k_{i}$ $(i=1,\ldots,n)$. The resulting model after this phase constitutes our final embedding model.
@@ 2.2.2 Expert Iteration
Another key component of REAL-Prover is the prover model within the tactic generator. To train a strong prover model, we combine expert iteration to collect synthetic data, as shown in Figure˜1(b). We begin by fine-tuning a prover model using open-source datasets. This initial model is then employed within REAL-Prover to solve formal statements produced by HERALD-AF, generating additional high-quality data. The newly collected data is incorporated into the supervised fine-tuning dataset to train an improved model. This refined model is then used to tackle challenging formal statements that the previous model could not solve. We update our model continuously in this iteration process.
@@ 3.1 Implementation Detail
We perform supervised fine-tuning on the base models Qwen2.5-Math-7B [yang2024qwen2] using a learning rate of $5\times 10^{-5}$ with a cosine decay scheduler and a maximum context length of 8192 tokens. Our training data comprises the following sources, totaling 210,420 state-tactic pairs:
  1. Post-processed formal proofs collected via expert iteration. Each proof is parsed into state-tactic pairs using Jixia, and the corresponding prompts are enriched with natural language descriptions, the proof context, the current proof state, and retrieval information from LeanSearch-PS. The dataset comprises 15,071 algebra problems from undergraduate and graduate mathematics texts, and 19,633 NuminaMath problems [numina_math_datasets], resulting in 25,818 and 30,589 state-tactic pairs, respectively, after post-processing.
@@
  4. The augmented Lean-Workbook dataset [ying2024leanworkbooklargescalelean]. We enhance this dataset by adding retrieval information from LeanSearch-PS to each example, while keeping the original structure intact. The resulting dataset comprises 43,192 state-tactic pairs.
@@
We develop two dense retrievers from E5-mistral-7b-instruct [wang2023improving] with Tevatron [Gao2022TevatronAE] training framework. For both initial training and hard negative enhanced training, we use LoRA to accelerate training, and set the learning rate as $2\times 10^{-5}$, the batch size as 128, the number of epoch as 1, and other hyperparameters as their default values. Queries are formal statements with a maximum length of 128 and passages are formal theorems with a maximum length of 256. For the hard negative examples, we first embed all statements and theorems with the initial trained embedding model, and then randomly select one passage from the top 30 to top 100 most similar ones for each query as its hard negative premise.
@@
In our experiments, considering the overall computational budget, we set the number of passes and samples per step to 64. We set our LLM’s temperature to 1.5. Additionally, we set the hyperparameter $\alpha$ of the score function to 0.5, consistent with the configuration used in BFS-Prover [xin2025bfsproverscalablebestfirsttree]